### v1

@@ 2.4 Unimodular regularization
In this subsection, we introduce Unimodular Regularization (UR), a novel technique developed as part of the MH setup. This method involves imposing specific constraints on the values of the latent space and its derivatives, enabling the MH framework to generalize effectively across variations within the family of DEs. By regularizing the latent space in this manner, we enhance its adaptability and ensure robust transfer learning capabilities. The method’s theoretical foundations and technical implementation are detailed below.
@@
During the training process, the body of the MH will learn a latent space $H_{i}(x^{\mu})$ that generalizes common features across different solutions to the DE within the same family. However, the choice of this set of functions is not unique, as the DEs are not exactly solved during training [^4] The DEs are solved by PINNs to the accuracy given by the square root of the loss function at a certain stage along the training., and the heads simultaneously learn how to combine these functions to produce specific solutions non-linearly. This introduces a potential problem: some learned latent spaces may not generalize well to regions of the parameter space outside the training range. Specifically, if the latent space exhibits rapid variation with respect to the input parameters $x^{\mu}$, TL to new parameter ranges may become problematic. In such cases, the frozen latent space may take significantly different values for variations within the family of DEs that are close to each other. This often happens when the equations are stiffer in these ranges [^5] This is typically the case in MH techniques, which are often used for TL in parameter ranges where the DEs are stiffer.. The head, in turn, would need to produce arbitrarily large responses to compensate for these rapid variations, making the solution unreliable.
To address this issue, we have developed a novel method called Unimodular Regularization (UR). This approach quantifies the values of the latent space and its derivatives by computing the associated metric tensor. By constraining the behavior of the latent space in this way, we ensure improved generalization and stability during TL, particularly in regions where the equations are more challenging.This is inspired by an alternative theory to General Relativity called Unimodular Gravity. In this theory the determinant of the metric tensor is picked to be $-1$ and the Gauge Group is reduced from diffeomorphisms (General Coordinate Transformations) to Transverse Diffeomorphisms (TDiff). These transformations are the ones that leave the determinant of the metric invariant. For more details see [28, 29].
We begin by embedding the latent space into the input space. This is achieved by parametrizing the latent space as a hyper-surface using the inputs $x^{\mu}$. To formalize this, we group all components into the following vectorial quantity
@@
where the total dimension of $\vec{\Omega}$ is given by $M=d+n$. Using standard techniques from differential geometry, we can compute the induced metric tensor on the hyper-surface with the following expression:
$g_{\mu\nu}\,=\,\frac{\partial\vec{\Omega}}{\partial x^{\mu}}\cdot\frac{\partial\vec{\Omega}}{\partial x^{\nu}}.$ (7)
Here, $\cdot$ denotes the usual scalar product [^6] The scalar product used here is defined as $\vec{x}\cdot\vec{y}=x_{1}y_{1}+x_{2}y_{2}+\ldots+x_{n}y_{n}$ where the dimension of each of the vectors is $n$.. It is important to note that $g_{\mu\nu}$ is a matrix of dimension $n\times n$. Notably, as the metric tensor $g_{\mu\nu}$ is computed according to (7), some entries of this matrix will be proportional to the derivatives of the latent space with respect to its inputs. These entries, therefore, quantify the sensitivity of $H_{i}(x^{\mu})$ to the inputs.
@@
The implementation of expression (7) is straightforward, as the required derivatives can be computed analytically through automatic differentiation. This results in an $n\times n$ matrix for each point in the batch, yielding a total size $=\text{batch size}\times n\times n$. To simplify this structure, we use expression (8) to compute a scalar quantity for each point in the batch, reducing the dimensions to match the batch size. This scalar function quantifies the sensitivity, or rate of change, of the latent space with respect to the chosen input points.
@@
The choice of using $\sqrt{g}$ is motivated by General Relativity. However, we expect to obtain similar results using $g$ instead of $\sqrt{g}$. Also note that because of the definition of the metric tensor, its determinant is always positive.
Since computing $g$ is computationally intensive, we only considered this additional loss every $N$ epochs. For epochs that are not divisible by $N$, the term is set to zero. Finally, if the model involves multiple bodies (as is often the case when the DE system includes more than one unknown function), the metric must be computed separately for each body. Consequently, for every unknown function in the system, a corresponding term of the form (9) is added to the total loss.
@@ 3.1 Flame equation
where we have added the $+1$ factor in the denominator to avoid dividing by zero.
@@
For training, we sampled 100 points in the $t$direction between the time bounds in (10), adding small random noise at each epoch to prevent overfitting. When training the MH body, we used $\delta=0.02$ (the smallest value in the range) to ensure the widest time intervals. For UR, we set $\lambda=5\cdot 10^{-7}$ and imposed the additional loss term every 100 epochs.
Figure 4: Training of the MH body for the flame equation. Top: Metric determinant without UR (left) and with UR (right). Note the difference in one to two orders of magnitude for the values of $g$ in each case. Bottom: Solutions to the flame equation $y^{\text{NN}}(t)$ (solid lines) compared with the numerical-RK4 solutions (dahsed lines) for each of the heads, for values of the parameter $\delta\in[0.02,0.04]$. As noted in the text, these solutions look identical for both cases with and without UR in the MH training process.
Figure 5: TL for the flame equation for $\delta=0.015$. Top: Loss functions for cases without UR (left) and with UR (right). TL starts at epoch $1.5\cdot 10^{6}$. Bottom left: Solutions for $y(t)$ from a body with UR (solid red line), without UR (solid blue), compared to the numerical solution (dashed green). Bottom right: Relative error (in %) for cases with UR (red) and without UR (blue), as defined in Eq. (11). In the legend, we show the value of the root mean square (rms) for each case.
@@ 3.2 Van der Pol oscillator
Figure 6: Training of the MH body for the van der Pol oscillator. Top: Metric determinant of the $y(t)$ body without UR (left) and with UR (right). Bottom: Solutions to the van der Pol oscillator $x^{\text{NN}}(t)$ (solid lines) compared with the numerical-RK4 solutions (dahsed lines) for each of the heads, for values of the parameter $a\in[0.0,1.5]$. The solutions look identical for both cases with and without UR in the MH training process.
Figure 7: TL of van der Pol oscillator for $a=1.75$. Top: Loss functions for the TL without UR (left) and with UR (right) on the MH bodies. TL starts at epoch $3\cdot 10^{6}$. Bottom left: Solution for $x^{\text{NN}}(t)$ for cases without UR (solid blue), with UR (solid red) compared to the RK4-numerical solution (dashed green). Bottom right: Relative error (in %) for cases with UR (red) and without UR (blue). In the legend, we show the value of the root mean square (rms) for each case.
@@
[…]
@@
$\frac{dx}{dt}\,&=\,\rho y\\ \frac{dy}{dt}\,&=\,\rho a(1-x^{2})+\rho x\end{aligned}\right.\qquad\text{IC: }\left\{\begin{aligned} x(0)\,&=\,1\\ y(0)\,&=\,0.\end{aligned}\right.$ (13)
@@ 3.3 Einstein Field Equations in AdS background
Moreover, its resolution involves solving a system of seven ordinary highly non-linear ODEs for approximately $70$ BCs, and recovering a free-form function appearing within the equations that is uniquely determined by these BCs. Consequently, this stands out as one of the most challenging setups for testing the algorithms introduced thus far.
@@
    Select a specific $V(\phi)$.  [^7] The function $V(\phi)$ must satisfy certain constraints on its value and derivatives near $\phi=0$ to ensure the resulting equations are well-defined. For a detailed discussion, see [22, 30]. A chosen $V(\phi)$ specifies an entire family of ODEs, with different BCs giving rise to variations within that family.
@@
Let us now discuss the inverse problem: we take a sample of $N$ points from a chosen $S(T)$ relation, which specifies a bundle $\{\text{BC}_{i}\}_{i=1}^{N}$. Consequently, the “full solution” $\{\vec{\psi}_{i}(u)\}_{i=1}^{N}$ becomes a set of $N$ different solutions to Eqs. (14a)–(14g) that must be found. Even though we know there is a corresponding function $V(\phi)$ related to the chosen $S(T)$/BCs, we do not actually know this function; it remains a free function determined by the solutions $\phi_{i}(u)\in\vec{\psi}_{i}(u)$ for each $i=1,\dots,N$.
@@
Subsequently, we tackle the inverse problem using selections of $S(T)$ with known morphological relationships to $\phi_{M}$. While a general choice of $S(T)$ might not be governed by a single parameter, within the same physical context, the correspondence between $S(T)$ and $V(\phi)$ remains intact. It is important to note that the link between scale separation and stiffness, while observed and utilized in our framework, is not derived from a formal mathematical proof but rather from empirical and heuristic understanding. Therefore, even though the stiffness of our Einstein Field Equations (EFE) problem is influenced by the morphology of $S(T)$, we refer to $\phi_{M}$ as a “stiffness parameter” analogous to the parameter $\delta$ in the flame equation (10) and the parameter $a$ in the Van der Pol (VdP) equation (12). Lower values of $\phi_{M}$ exacerbate the problem’s stiffness, making it increasingly difficult to recover the free function. This difficulty persists until the original PINN setup either fails to recover $V(\phi)$ within an acceptable error margin or, in cases of extreme stiffness, fails to recover it altogether, including its most fundamental features, such as the position $\phi_{H}$ and scale $V(\phi_{H})$ of the critical points. For a more in-depth discussion on the stiffness of the problem, the features in $V(\phi)$, and the morphologies of $S(T)$, we refer the reader to our previous work [22], particularly Figures 1, 2, and 13.
@@
Figure 8: Training of the MH body for the EFE. Top: Metric determinant for cases without UR (left) and with UR (right), for the head corresponding to the parameter $\phi_{M}=1$, for one of the solution functions, $\nu_{\Sigma}(u)\in\vec{\psi}(u)$. In both panels, the black, dashed, horizontal line corresponds to the ODE problem given by the BC that is marked with a purple star in the bottom, left panel (theory with $\phi_{M}=1$). Bottom left: Different input points given to the body that come from different choices of $S(T)_{\phi_{M}}$. Marked with a star, we show the BC point with index $i=10$ for all $\phi_{M}$. This point approximately tracks where the equations become stiffer within the corresponding $\phi_{M}$ family. Bottom right: The recovered $V^{\text{NN}}(\phi;\phi_{M})$ coming from the MH body are shown in solid lines, while the theoretical $V^{\text{th}}(\phi;\phi_{M})$ are shown in dashed lines.
As explained in Sec. 2.2 and 2.3, we then use the frozen latent space to perform TL to stiffer regimes. In our approach, we do so for two new $S(T)$ curves corresponding to $\phi_{M}=0.9$ and $\phi_{M}=0.7$. We test UR on both cases and display our findings in figures 8 (training of the body), 9 (case with $\phi_{M}=0.9$) and 10 (case with $\phi_{M}=0.7$).
Before we present the results, we summarize the architecture and training procedure for the EFE case. These procedures are identical in scenarios both with and without UR to ensure a fair comparison of their performance. Each solution function in $\vec{\psi}(u)=~{}\left(\tilde{\Sigma}(u),\tilde{A}(u),\phi(u),\nu_{\Sigma}(u),\nu_{A}(u),\nu_{\phi}(u)\right)$ is produced by an individual FCNN. The architecture of these NNs consists of $[32,32,32,32,128]$ layers for the main bodies and $[16,16]$ for the five heads, all utilizing Tanh activation. The free function $V^{\text{PINN}}(\phi)$ is derived from another FCNN with a body structured as $[32,32,32,128]$ and five heads composed of $[64,64]$, with SiLU as the activation for both the body and heads. The optimizer used for body training and transfer learning is adam, starting with an initial learning rate of $1\cdot 10^{-3}$. As part of the scheduling process, the learning rate is reduced by $1.5\,\%$ every $5000$ epochs, both in training the body and in transfer learning. In scenarios with UR, it is imposed in the loss calculation every 500 iterations, with a global coefficient $\lambda=5\cdot 10^{-8}$.
TL under UR exhibited significant improvement in recovering the free function $V(\phi;\phi_{M})$ and the associated $S(T)_{\phi_{M}}$. Specifically, for TL to $\phi_{M}=0.9$, our UR approach accomplished a twofold improvement in the recovery of $V(\phi,0.9)$, reducing root mean square (rms) from roughly $5\%$ to about $3\%$. Consequently, this also signified a twofold improvement in $S(T)_{0.9}$, with rms decreasing from around $13\%$ to about $6\%$. In the context of $\phi_{M}=0.7$, results are less precise overall due to significant increases in problem stiffness. Nevertheless, there is a threefold enhancement through UR in retrieving $V(\phi,0.7)$, with rms dropping from approximately $64\%$ to approximately $22\%$, and a sevenfold improvement in $S(T)_{0.7}$, reducing rms from around $178\%$ to roughly $26\%$. In both scenarios, UR demonstrates improvements in recovery, and for $\phi_{M}=0.7$, it enables a closer TL approximation, whereas the conventional MH case resulted in extraordinarily large errors.
Figure 9: TL of EFE for $\phi_{M}=0.9$. Top left: Recovered functions $V^{\text{PINN}}(\phi)$ through TL without UR (solid blue), and with UR (solid red), compared with the theoretical function (dashed green). Top right: $S^{\text{PINN}}(T)$ relations associated to the recovered $V^{\text{PINN}}(\phi)$ without and with UR, compared to the theoretical one (which is also the initial input in the TL procedure). Bottom: Relative errors for both the recovered $V^{\text{PINN}}(\phi)$ (left) and associated $S^{\text{PINN}}(T)$ (right) for the cases without UR (blue) and with UR (red). In the legend, we show the value of the root mean square (rms) for each case.
Figure 10: TL of EFE for $\phi_{M}=0.9$. Top left: Recovered functions $V^{\text{PINN}}(\phi)$ through TL without UR (solid blue), and with UR (solid red), compared with the theoretical function (dashed green). Top right: $S^{\text{PINN}}(T)$ relations associated to the recovered $V^{\text{PINN}}(\phi)$ without and with UR, compared to the theoretical one. Bottom: Relative errors for both the recovered $V^{\text{PINN}}(\phi)$ (left) and associated $S^{\text{PINN}}(T)$ (right) for the cases without UR (blue) and with UR (red). In the legend, we show the value of the root mean square (rms) for each case.

### v2

@@ 2.5 Unimodular regularization
In this subsection, we introduce Unimodular Regularization (UR), a technique developed as part of the MH setup. This method involves imposing specific constraints on the values of the latent space and its derivatives, enabling the MH framework to generalize effectively across variations within the family of DEs. By regularizing the latent space in this manner, we enhance its adaptability and ensure robust transfer learning capabilities. The method’s theoretical foundations and technical implementation are detailed below.
@@
During the training process, the body of the MH will learn a latent space $H_{i}(x^{\mu})$ that generalizes common features across different solutions to the DE within the same family. However, the choice of this set of functions is not unique, as the DEs are not exactly solved during training (the DEs are solved by PINNs to the accuracy given by the square root of the loss function at a certain stage along the training), and the heads simultaneously learn how to combine these functions to produce specific solutions non-linearly. This introduces a potential problem: some learned latent spaces may not generalize well to regions of the parameter space outside the training range. Specifically, if the latent space exhibits rapid variation with respect to the input parameters $x^{\mu}$, TL to new parameter ranges may become problematic. In such cases, the frozen latent space may take significantly different values for variations within the family of DEs that are close to each other. This often happens when the equations are stiffer in these ranges (this is typically the case in MH techniques, which are often used for TL in parameter ranges where the DEs are stiffer). The head, in turn, would need to produce arbitrarily large responses to compensate for these rapid variations, making the solution unreliable.
To address this issue, we have developed a method called Unimodular Regularization (UR). This approach quantifies the values of the latent space and its derivatives by computing the associated metric tensor. By constraining the behavior of the latent space in this way, we ensure improved generalization and stability during TL, particularly in regions where the equations are more challenging. This is inspired by an alternative theory to General Relativity called Unimodular Gravity. In this theory the determinant of the metric tensor is picked to be 1-1 and the Gauge Group is reduced from diffeomorphisms (General Coordinate Transformations) to Transverse Diffeomorphisms (TDiff). These transformations are the ones that leave the determinant of the metric invariant. For more details see [36, 37].
We begin by using inputs $x^{\mu}$ as coordinates of a higher-dimensional manifold $\vec{\Omega}$ that also contains the latent space $H_{i}(x^{\mu})$. This is achieved by parametrizing the latent space as a hyper-surface of the inputs $x^{\mu}$. To formalize this, we group all components into the following quantity
@@
where $i=1,\dots,d$, and $\mu=1,\dots,n$. The total dimension of $\vec{\Omega}$ is given by $M=d+n$. Concretely, $H_{i}(x^{\mu})$ is a $d$-dimensional hyper-surface parametrized by $n$ variables, embedded in an $M$-dimensional space. Using standard techniques from differential geometry, we can compute the induced metric tensor on the hyper-surface with the following expression:
$g_{\mu\nu}\,=\,\frac{\partial\vec{\Omega}}{\partial x^{\mu}}\cdot\frac{\partial\vec{\Omega}}{\partial x^{\nu}}.\vskip 8.5359pt$ (7)
Here, $\cdot$ denotes the usual scalar product, defined as $\vec{x}\cdot\vec{y}=x_{1}y_{1}+x_{2}y_{2}+\ldots+x_{n}y_{n}$ where the dimension of each of the vectors is $n$. It is important to note that $g_{\mu\nu}$ is a matrix of dimension $n\times n$. Notably, as the metric tensor $g_{\mu\nu}$ is computed according to (7), some entries of this matrix will be proportional to the derivatives of the latent space with respect to its inputs. These entries, therefore, quantify the sensitivity of $H_{i}(x^{\mu})$ to the inputs.
@@
The implementation of equation (7) is relatively straightforward in the PINNs setup. In order to compute the metric tensor it is necessary to compute the derivatives of the latent space with respect to the inputs. These are computed analytically using automatic differentiation. This algorithm is already implemented on most of machine learning pre-built packages. For the details of the implementation we refer the reader to our GitHub repository where all the codes can be found (see the “Data and Code Availability” section). This results in an $n\times n$ matrix for each point in the batch, yielding a total size $\text{batch size}\times n\times n$. To simplify this structure, we use expression (8) to compute a scalar quantity for each point in the batch, reducing the dimensions to match the batch size. This scalar function quantifies the sensitivity, or rate of change, of the latent space with respect to the chosen input points.
@@
Although this regularization seems similar to Jacobian regularization (JR) technique [38], it is fundamentally different in its geometrical interpretation; the determinant of the metric $g$ has a clear meaning in terms of differential geometry as the volume of a differential of the hyper-volume in which $H_{i}(x^{\mu})$ lives. The pseudo-Jacobian present in JR lacks this interpretation. Also, we find that JR as a regularizing technique is implicit when performing UR. A detailed explanation can be found in the “Supplementary Note 1”.
A qualitative explanation of the effect of adding the UR loss into the loss landscape is the following. We know that there must exist some combination of parameters (some state of the NNs) such that the NN output functions are a solution of the DEs up to some precession. This approximate solution corresponds to some local minimum in the high-dimensional space of parameters in which the loss landscape lives. However the “throat” that will bring the NN’s state into this minimum is very narrow. We hypothesize that adding the UR loss (9) has the effect of widening this throat. Then, it will be more likely for the NN to fall into the local minimum during the TL phase. This hypothesis is compatible with what we have observed in our tests, and with the losses shown in the figures of the following subsections.
The choice of using $\sqrt{g}$ is motivated by General Relativity. However, we expect to obtain similar results using $g$ instead of $\sqrt{g}$. Also note that because of the definition of the metric tensor, its determinant is always positive. Moreover, using this object rather than any other differential geometry quantities presents various advantages, among which we underline the following
  • It is the simplest quantity one can compute. The determinant of the metric $g$ will only contain up to first derivatives of the latent space with respect to the inputs. Any other geometrical scalar of the manifold (for example, Ricci scalar) will imply computing second order derivatives of the latent space. During the tests, we found out that computing this quantity once can take up to $30$ minutes of computing time (this quantity needs to be computed recursively, which makes it unfeasible).
  • The geometrical interpretability of this quantity is also clear: it measures the differential of volume of the latent space as a hyper-volume living in a higher-dimensional manifold parametrized by the inputs as coordinates. Consequently, the larger it is, the more latent space volume (or information) is captured within a coordinate differential.
The interested reader can find a brief analysis of the relation of UR with the Lipschitz constant in the “Supplementary Note 2”, as further motivation for the use of the metric determinant $\sqrt{g}$ in our regularizing technique.
Since computing $g$ is computationally intensive, we only considered this additional loss every $N$ epochs. For epochs that are not divisible by $N$, the term is set to zero. It is important for the stability of the model during training, specially when doing UR, to keep both the learning rate and the $\lambda$ parameter sufficiently small. The hyperparameter $N$ needs also to be chosen properly: if NN is too small, computational time increases, also the NN reduces its flexibility to fit solutions. If NN is too large, the regularization of the latent space is lost to all the training epochs without it. A proper $N$ will yield regularization that facilitates TL while keeping enough freedom of the NN to fit solutions. If the learning rate is very large the network will be “pushed out” of the local minima when the loss (9) is taken into account. We have found both the values for $\lambda$ and the learning rate by performing different tests of the algorithm. It is important to note that these values will change depending on the ODEs that one wants to solve. The specific values for each case are indicated in the “Results” section. Finally, if the model involves multiple bodies (as is often the case when the DE system includes more than one unknown function), the metric must be computed separately for each body. Consequently, for every unknown function in the system, a corresponding term of the form (9) is added to the total loss.
The overall scaling of computational time of the problem given UR is as follows: since we need to compute the derivatives in (7) through AD, training time scales as $b\times n\times d$ (being $b$ the batch size, $n$ number of inputs $x^{\mu}$, and $d$ dimension of the latent space). It scales linearly with the number of heads (each iteration implies a forward pass and a backward pass per head). It also scales linearly with the number of unknown functions in the DEs. It is complicated to give a quantitative statement of the scaling with the stiffness parameter, given that there is not a single definition of stiffness as described in the “Introduction” section. Qualitatively, as stiffness of the problem increases,the number of iterations also increases both for the MH body training, and the TL phase.
@@ 3.2 Flame equation
where we have added the +1+1 factor in the denominator to avoid dividing by zero.
@@
For training, we sampled 100 points in the $t$ -direction between the time bounds in (10), adding small random noise at each epoch to prevent overfitting. When training the MH body, we used $\delta=0.02$ (the smallest value in the range) to ensure the widest time intervals. For UR, we set $\lambda=5\cdot 10^{-7}$ and imposed the additional loss term every 100 epochs.
Figure 4: Training of the Multi-Head body for the flame equation. $y^{\text{NN}}(t)$ is the ODE solution with independent variable $t$ given by the NN. $y^{RK54}(t)$ is the ODE numerical solution. $\delta$ is the initial condition on $y(t)$. $g$ is the value of the determinant of the metric. (a): Metric determinant without UR. (b): Metric determinant with UR. Note the difference in one to two orders of magnitude for the values of $g$ in each case. (c): Solutions to the flame equation $y^{\text{NN}}(t)$ (solid lines) compared with the numerical-RK4 solutions (dahsed lines) for each of the heads, for values of the parameter labeling different ICs, $y(0)=\delta\in[0.02,0.04]$. As noted in the text, these solutions look identical for both cases with and without UR in the MH training process.
Figure 5: Transfer learning for the flame equation. $y(t)$ is the solution to the ODE, for numerical (RK4), and for NN with ($y_{\text{UR}}^{\text{NN}}$) and without UR ($y^{\text{NN}}$). Relative error (RE) between the numerical and the NN solution. The initial condition is $y(0)=\delta=0.015$. (a): Loss functions for cases without UR. (b): Loss functions for cases with UR. Transfer learning starts at epoch $1.5\cdot 10^{6}$. (c): Solutions for $y(t)$ from a body with UR (solid red line), without UR (solid blue), compared to the numerical solution (dashed green). (d): RE (in %) for cases with UR (red) and without UR (blue), as defined in Eq. (11). In the legend, we show the value of the root mean square (rms) for each case.
@@ 3.3 Van der Pol oscillator
Figure 6: Training of the multi-head body for the van der Pol oscillator. $x(t)$ is the solution to the ODE numerical (RK4) and given by the NN. $a$ is the stiffness parameter appearing in the ODE and $g$ is the determinant of the metric. (a): metric determinant of the $y(t)$ body without UR. (b): metric determinant of the $y(t)$ body with UR. (c): solutions to the van der Pol oscillator $x^{\text{NN}}(t)$ (solid lines) compared with the numerical-RK4 solutions (dashed lines) for each of the heads, for values of the parameter $a\in[0.0,1.5]$. The solutions look identical for both cases with and without UR in the MH training process.
Figure 7: Transfer learning of van der Pol oscillator. $x(t)$ is the numerical (RK4) and NN solution to the ODE. The stiffness parameter appearing in the ODE, for this case of TL $a=1.75$. RE is the relative erorr between the numerical the NN solutions. (a): Loss functions for the TL without UR. (b): Loss functions for the TL with UR. TL starts at epoch $3\cdot 10^{6}$. (c): NN-solution $x^{\text{NN}}(t)$ for cases without UR (solid blue), with UR (solid red) compared to the RK4-numerical solution (dashed green). (d): Relative error (in %) for cases with UR (red) and without UR (blue). In the legend, we show the value of the root mean square (rms) for each case.
@@
$\frac{d^{2}x}{dt^{2}}-a(1-x^{2})\frac{dx}{dt}+x\,=\,0;\qquad x(0)\,=\,1;\qquad\left.\frac{dx}{dt}\right|_{t=0}\,=\,0,$ (12)
@@
$\left\{\begin{aligned} \frac{dx}{dt}\,&=\,\rho y\\ \frac{dy}{dt}\,&=\,\rho a(1-x^{2})+\rho x\end{aligned}\right.\qquad\text{IC: }\left\{\begin{aligned} x(0)\,&=\,1\\ y(0)\,&=\,0.\end{aligned}\right$. (13)
@@ 3.4 Einstein Field Equations in AdS background
Moreover, its resolution involves solving a system of seven ordinary highly non-linear ODEs for approximately $70$ BCs, and recovering a free-form function appearing within the equations that is uniquely determined by these BCs. Consequently, this stands out as one of the most challenging setups for testing the algorithms introduced so far.
@@
    Select a specific $V(\phi)$: the function $V(\phi)$ must satisfy certain constraints on its value and derivatives near $\phi=0$ to ensure the resulting equations are well-defined. For a detailed discussion, see [39, 40]. A chosen $V(\phi)$ specifies an entire family of ODEs, with different BCs giving rise to variations within that family.
@@
Let us now discuss the inverse problem: we take a sample of $N$ points from a chosen $S(T)$ relation, which specifies a bundle $\{\text{BC}_{i}\}_{i=1}^{N}$. Consequently, the “full solution” $\{\vec{\psi}_{i}(u)\}_{i=1}^{N}$ becomes a set of $N$ different solutions to Eqs. (14a)–(14g) that must be found. Even though we know there is a corresponding function $V(\phi)$ related to the chosen S(T)S(T)/BCs, we do not actually know this function; it remains a free function determined by the solutions $\phi_{i}(u)\in\vec{\psi}_{i}(u)$ for each $i=1,\dots,N$.
@@
Subsequently, we tackle the inverse problem using selections of $S(T)$ with known morphological relationships to $\phi_{M}$. While a general choice of $S(T)$ might not be governed by a single parameter, within the same physical context, the correspondence between $S(T)$ and $V(\phi)$ remains intact. It is important to note that the link between scale separation and stiffness, while observed and utilized in our framework, is not derived from a formal mathematical proof but rather from empirical and heuristic understanding. Therefore, even though the stiffness of our EFE problem is influenced by the morphology of $S(T)$, we refer to $\phi_{M}$ as a “stiffness parameter” analogous to the parameter $\delta$ in the flame equation (10) and the parameter $a$ in the Van der Pol (VdP) equation (12). Lower values of $\phi_{M}$ exacerbate the problem’s stiffness, making it increasingly difficult to recover the free function. This difficulty persists until the original PINN setup either fails to recover $V(\phi)$ within an acceptable error margin or, in cases of extreme stiffness, fails to recover it altogether, including its most fundamental features, such as the position $\phi_{H}$ and scale $V(\phi_{H})$ of the critical points. For a more in-depth discussion on the stiffness of the problem, the features in $V(\phi)$, and the morphologies of $S(T)$, we refer the reader to Figures 1, 2, and 13 of [39].
In an attempt to invert this problem given data $S(T)$, one might be concerned that the sampled data does not cover the whole observable space, and thus the inverted function $V(\phi)$ cannot be fully recovered. However, this is not the case in this setup. The “observed data” $S(T)$ are in fact synthetic, obtained by numerically solving the direct EFE problem for a potential $V$ drawn from the same family $V(\phi,\phi_{M})$ at a fixed $\phi_{M}$. Each sample in $S(T)$ corresponds to one set of BCs, fully defining the solution and thus restricting the solution space to a single solution. We further truncate $S(T)$ at very high temperatures—an approximation justified by the problem’s symmetry, since beyond the asymptotic regime ($S\propto T^{3}$; see [39]), no new information is gained. Moreover, the sampled data for the independent variable $u$ will cover most of the observable space. It ranges from 0 to $1$ which in the physical context correspond to the whole size of the spacetime in which EFEs are being solved. It is always sampled using a Chebyshev distribution with noise. Thus, we expect that this range will be entirely covered when the model is trained for a sufficiently large number of iterations.
In this context, we introduce a MH approach: we supply our MH PINN framework with five increasingly stiff curves $S(T)_{\phi_{M}}$, corresponding to $\phi_{M}=\{3.0,\,2.0,\,1.5,\,1.08,\,1.0\}$ (see Fig. 8, bottom panels). Through its shared body, the MH PINN learns the latent space of solutions and free functions associated with the various morphologies of $S(T)_{\phi_{M}}$. Subsequently, each head projects down from this shared latent space to produce a specific $V^{\text{PINN}}(\phi;\phi_{M})$, as well as the corresponding ODE solutions $\vec{\psi}_{i,\phi_{M}}$, for the BCs specified by each $\{T_{i},S_{i}\}_{\phi_{M}}$.
Figure 8: Training of the multi-head body for the Einstein Field Equations. g is the metric determinant. $\phi_{M}$ is the stiffness parameter. $\phi$ is one of the unknown solution functions, and $V(\phi)$ the unknown potential. $\Sigma_{h}$ and $V_{a,h}$ are different boundary conditions of the unknown functions. (a) Metric determinant for cases without UR. (b) Metric determinant for cases without UR. The head corresponds to the parameter $\phi_{M}=1$, for one of the solution functions, $\nu_{\Sigma}(u)\in\vec{\psi}(u)$. In both panels, the black, dashed, horizontal line corresponds to the ODE problem given by the BC that is marked with a purple star in panel (c) (theory with $\phi_{M}=1$). (c): Different input points given to the body that come from different choices of $S(T)_{\phi_{M}}$. Marked with a star, we show the BC point with index $i=10$ for all $\phi_{M}$. This point approximately tracks where the equations become stiffer within the corresponding $\phi_{M}$ family. (d): Recovered NN potentials $V^{\text{NN}}(\phi;\phi_{M})$ coming from the MH body are shown in solid lines, while the theoretical $V^{\text{th}}(\phi;\phi_{M})$ are shown in dashed lines.
As explained in the “Multihead” and the “Training and Transfering” subsections, we then use the frozen latent space to perform TL to stiffer regimes. In our approach, we do so for two new $S(T)$ curves corresponding to $\phi_{M}=0.9$ and $\phi_{M}=0.7$. We test UR on both cases and display our findings in figures 8 (training of the body), 9 (case with $\phi_{M}=0.9$) and 10 (case with $\phi_{M}=0.7$).
Before we present the results, we summarize the architecture and training procedure for the EFE case. These procedures are identical in scenarios both with and without UR to ensure a fair comparison of their performance. Each solution function in $\vec{\psi}(u)=~\left(\tilde{\Sigma}(u),\tilde{A}(u),\phi(u),\nu_{\Sigma}(u),\nu_{A}(u),\nu_{\phi}(u)\right)$ is produced by an individual FCNN. The architecture of these NNs consists of $[32,32,32,32,128]$ layers for the main bodies and $[16,16]$ for the five heads, all utilizing Tanh activation. The free function $V^{\text{PINN}}(\phi)$ is derived from another FCNN with a body structured as $[32,32,32,128]$ and five heads composed of $[64,64]$, with SiLU as the activation for both the body and heads. The optimizer used for body training and TL is adam, starting with an initial learning rate of $1\cdot 10^{-3}$. As part of the scheduling process, the learning rate is reduced by $1.5\,\%$ every $5000$ epochs, both in training the body and in TL. In scenarios with UR, it is imposed in the loss calculation every 500 iterations, with a global coefficient $\lambda=5\cdot 10^{-8}$.
TL under UR exhibited significant improvement in recovering the free function $V(\phi;\phi_{M})$ and the associated $S(T)_{\phi_{M}}$. Specifically, for TL to $\phi_{M}=0.9$ (Fig. 9), our UR approach accomplished a twofold improvement in the recovery of $V(\phi,0.9)$, reducing root mean square (rms) from roughly $5\%$ to about $3\%$. Consequently, this also signified a twofold improvement in $S(T)_{0.9}$, with rms decreasing from around $13\%$ to about $6\%$. In the context of $\phi_{M}=0.7$, results are less precise overall due to significant increases in problem stiffness. Nevertheless, there is a threefold enhancement through UR in retrieving $V(\phi,0.7)$, with rms dropping from approximately $64\%$ to approximately $22\%$, and a sevenfold improvement in $S(T)_{0.7}$, reducing rms from around $178\%$ to roughly $26\%$ (Fig. 10). In both scenarios, UR demonstrates improvements in recovery, and for $\phi_{M}=0.7$, it enables a closer TL approximation, whereas the conventional MH case resulted in extraordinarily large errors.
Figure 9: Transfer learning of Einstein Field Equations for a non-pronounced first order phase transition. $V(\phi)$ is the free form potential recovered by the NN (PINN) and the theoretical one (th). $S(T)$ is the thermodynamical curve (entropy as a function of temperature) recovered by the NN (PINN) and the theoretical one (th). RE refers to relative error between the NN model and the theoretical result. (a) Recovered functions $V^{\text{PINN}}(\phi)$ through TL without UR (solid blue), and with UR (solid red), compared with the theoretical function (dashed green), corresponding to $\phi_{M}=0.9$. (b) $S^{\text{PINN}}(T)$ relations associated to the recovered $V^{\text{PINN}}(\phi)$ without and with UR, compared to the theoretical one (which is also the initial input in the TL procedure). (c): REs for the recovered $V^{\text{PINN}}(\phi)$ for the cases without UR (blue) and with UR (red). (d): RE for the associated $T^{\text{PINN}}(S)$ for the cases without UR (blue) and with UR (red). In the legend, we show the value of the root mean square (rms) for each case.